OpenStudy (anonymous):

Given the enthalpies of reaction below, determine the enthalpy change for the reaction: N2O (g) + NO2 (g) arrow 3NO (g). N2 (g) + O2 (g) arrow2NO (g) ΔH = +180.7 kJ 2NO (g) + O2 (g) arrow2NO2 (g) ΔH = -113.1 kJ 2N2O (g) arrow2N2 (g) + O2 (g) ΔH = -163.2 kJ Answer 155.6 kJ 130.6 kJ 311.3 kJ 65.3 kJ

OpenStudy (anonymous):

@ganeshie8 @texaschic101 @nincompoop

OpenStudy (anonymous):

@Somy

OpenStudy (anonymous):

SOMY!! Thanks the Gods, please help.

OpenStudy (somy):

hmm well firstly change these reactions to the way you want them to be

OpenStudy (anonymous):

yeah...how do I do that?

OpenStudy (somy):

N2 (g) + O2 (g) arrow2NO (g) ΔH = +180.7 kJ alright so here we have enthalpy change of formation of NO in the main reaction its 3 but here its 2 (coefficient meaning) i believe you can just do it like this +180.7 = 2 X= 3 X= +180.7*3/2 X= new enthalpy change to get 3 of NO [not sure tho]

OpenStudy (somy):

2NO (g) + O2 (g) arrow2NO2 (g) ΔH = -113.1 kJ alright here enthaply change is given for 2 of NO2 while in main reaction we have only 1 NO2 so if -113.1 kJ = 2 X= 1 X= -113.1/2 X= new enthalpy change of formation of 1 NO2

OpenStudy (anonymous):

Nm, I think I got the answer can you help with this one tho? hich of the following is an example of objects in thermal equilibrium? Answer ice warms in a glass before you can pour soda into the glass a spoon in a cup of hot soup absorbs heat a pan is placed on a stove burner and becomes hot two panes of glass are at the same temperature as air trapped between them

OpenStudy (anonymous):

I think it's d

OpenStudy (somy):

2N2O (g) arrow2N2 (g) + O2 (g) ΔH = -163.2 kJ in this reaction its rather formation of 2N2 (g) + O2 (g) but we need the opposite so if for forming 2N2 (g) + O2 (g) we enthalpy change is -163.2 kJ that means if we want the opposite effect = we can just change the sign to +163.2 kJ now the reaction looks like this 2N2 (g) + O2 (g)-> 2N2O (g) ΔH = +163.2 kJ again in the main reaction there are only 1 of N2O but here there are 2 of N2O so +163.2 kJ = 2 X= 1 X= +163.2 kJ/2 X= new enthalpy change of formation of N2O (g)

OpenStudy (somy):

so now you have everything so u can use the formula of Enthalpy change of reaction Enthalpy change of reaction = Enthalpy change of reactants- Enthalpy change of products

OpenStudy (somy):

do u think u can do it now?

OpenStudy (somy):

btw im not sure about the other question sorry

OpenStudy (anonymous):

Wow...thanks a bunch! :D

OpenStudy (somy):

you are welcome (๑>ᴗ<๑) :)))